OpenStudy (anonymous):

Write the equation for a circle that satisfies the given conditions. 1. Center (-4,-2) and the radius 14 units. 2.Center (5,-4) and the radius is 4 units. 3. Center (-8,7) and radius 1/2 4.Write an equation for the circle that satisfies the given set of conditions endpoints of a diameter at (11,18) and (-13,-19) 4.Write an equation for the circle the satisfies the given set of conditions. Center (8,-9) and passes through (21,22) 5.Write an equation for the circle the satisfies the given set of conditions. Center (4,-7) and the radius 5 units.

OpenStudy (zzr0ck3r):

the equation of a circle with center \((a,b)\) and radius \(r\) is given by \((x-a)^2+(y-b)^2=r^2\)

OpenStudy (anonymous):

i know its a lot and i am sorry!!

OpenStudy (zzr0ck3r):

if you are given center \((a,b)\) and a point it passes through \((c,d)\) then use the distance formula \(D=r=\sqrt{(a-c)^2+(b-d)^2}\) to get the radius and then use the equation above.

OpenStudy (anonymous):

use this eq.. (x-a)^2+(y-b)^2=r^2 here (a,b) are oordinates of point given and r is radius..

OpenStudy (anonymous):

Can yall please give me the answers, I have to get pass this test or my mom is going to strangle me. not really,but close. lol

OpenStudy (zzr0ck3r):

read the rules, we are not here to take your test for u.

OpenStudy (anonymous):

Whatever.